Find the formula of the compound with the percent composition shown

O, 41.26

P, 19.97

Ca 38.76


1
Expert's answer
2022-05-22T22:59:03-0400

1. We divide the mass fraction of each element by its relative atomic mass:

O — 41,26 / 16 = 2,57875

P — 19,97 / 31 = 0,6442

Ca — 38,78 / 40 = 0,9695

2. We divide each of the results into the smallest of them.(the smallest here is 0,6442):

O — 2,57875 / 0,6442 = 4

P — 0,6442 / 0,6442 = 1

Ca — 0,9695 / 0,6442 = 1,5

3. we can express the result in whole numbers. to do this, we multiply each of the results by 2 or 3, and so on. (until an integer is formed):

O — 4 * 2 = 8

P — 1 * 2= 2

Ca — 1,5 * 2 = 3

The result Ca = 3, P = 2, O = 8

The substance with this composition is Ca3(PO4)2 calcium phosphate

ANSWER: Ca3(PO4)2